OpenStudy (anonymous):

1) 4 - t = 3(t - 1)-5 a. 3 b. 6 c. no solution d. identity

OpenStudy (anonymous):

4 - t = 3(t - 1)-5 4 - t = 3t - 3 - 5 4 - t = 3t - 8 4+8 = 3t + t now find t

OpenStudy (anonymous):

first do pemdas